Requirement 1. Fil in the missing amounts. Begin by completing the income statement. Toad Company ncome Statement For the Year Ended December 31, 2012 Sales Cost of goods sold . Gross profit on sales.. . Administrative expenses .. Operating income 725,000 320,000 405,000 45,000 360,000 8,000 123,200 228,800 Net incomeNow fill in the missing amounts on the balance sheet. Toad Company Balance Sheet At December 31, 2012 Accounts payable.$ 13,500 9,100 2,000 Note payable 47,000 974,000 Contributed capital . . Accounts receivable . Equipment.... 560,000 Retained earnings . . $ 1,123,000Need help solving the Balance Sheet.
Answer #1

The key to solving the balance sheet lies in understanding the fact that the left-hand side and right-hand side of the balance sheet have equal total values. This essentially implies that Total Assets (Left-Hand Side) = Total Liabilities + Shareholder's Equity (Right-Hand Side).

Total Assets = Total Liabilities + Shareholder's Equity = $ 1123000

Total LIabilities + Shareholder's Equity = 1123000 = Retained Earnings + Contributed Capital + Note Payable + Accounts Payable

1123000 = Retained Earnings + 560000 + 9100 + 13500

Retained Earnings = 1123000 - 560000 - 13500 - 9100 = $ 540400

Cash + Accounts Receivable + Inventory + Equipment = 1123000

Cash + 12000 + 47000 + 974000 = 1123000

Cash = 1123000 - 974000 - 47000 - 12000 = $ 90000